# Homogeneous potentials, Lagrange’s identity and Poisson geometry

### v3

@@ 1 Introduction
$H=T+V=\frac{1}{2m_{1}}p_{1}^{2}+\cdots+\frac{1}{2m_{n}}p_{n}^{2}+V(q)\,,\qquad m_{i}>0\,$,
is the sum of kinetic energy $T$ and a smooth homogeneous function $V(q)$ of degree $k$, then Lagrange’s identity holds
@@
Jacobi used this identity considering the stability of the motion of systems of $n$ interacting particles with masses $m_{1},\ldots,m_{n}$ arising in celestial mechanics [1], see also [2]. Various generalizations and applications of Lagrange’s identity are discussed in the works of V.V. Kozlov [3, 4].
@@
In this note we prove that if potential $V(q)$ is a smooth homogeneous function of degree $k$, then for the Hamiltonian system (1.1,1.5) not only Lagrange’s identity (1.2) holds, but also exist an invariant Poisson bivector $\hat{P}$ and an invariant symplectic form $\hat{\omega}=\hat{P}^{-1}$, which cannot be obtained by differentiation and tensor operations from the basic tensor invariants $X,H,P$ and $\omega$.
@@ 1.1 Basic tensor invariants of Hamiltonian systems
$\Omega=\wedge^{n}\omega=dq_{1}\wedge\cdots\wedge dq_{n}\wedge dp_{1}\cdots\wedge dp_{n}\,,\qquad\mathcal{L}_{X}\Omega=0\,.$
@@
1.2 Invariant Volterra brackets
In 1895-1899 Vito Volterra proposed a formal Hamiltonian description of second-order homogeneous vector fields with two quadratic integrals of motion. Our aim is to present similar formal Hamiltonian description for homogeneous vector fields with one homogeneous integral of motion.
In [15] Volterra consider dynamical system
$\dot{x}_{k}=\sum a_{k}^{ij}x_{i}x_{j}\,,\qquad a_{k}^{ij}\in\mathbb{R}\,.$ (1.12)
with two quadratic first integrals
$F_{A}=\frac{1}{2}(Ax,x),\qquad F_{B}=\frac{1}{2}(Bx,x).$
If quadratic form $F_{A}$ is positive-definite and $\lambda_{1},\cdots,\lambda_{n}$ are pairwise different real roots of the characteristic polynomial
$\mbox{det}(B-\lambda A)=0,$
then a linear change of variables reduces the system of differential equations (1.12) to the form
$\dot{x}_{k}=\sum_{i<j\,i,j\neq k}c^{k}_{ij}(\lambda_{i}-\lambda_{j})x_{i}x_{j}$ (1.13)
where
$c^{j}_{ki}=c^{i}_{jk}=c^{k}_{ij}=-c^{k}_{ik}=-c^{i}_{kj}=-c^{i}_{jk},$
and other $c^{k}_{ij}$ are equal to zero. The Volterra proof is based on reduction of the pair of quadratic forms $F_{A}$ and $F_{B}$ to canonical form
$F_{A}=\frac{1}{2}\left(x^{2}_{1}+\cdots+x^{2}_{n}\right),\qquad F_{B}=\frac{1}{2}\left(\lambda_{1}x^{2}_{1}+\cdots+\lambda_{n}x^{2}_{n}\right).$
Then Volterra represents equations (1.13) by using the Jacobian of the first integrals
$\dot{x}_{k}=\sum_{i<j,\,i,j\neq k}c^{k}_{ij}\frac{\partial(F_{A},F_{B})}{\partial(x_{i},x_{j})}$
and by using almost Poisson bivectors
$X=P_{A}dF_{B}=P_{B}dF_{A}\,.$ (1.14)
Here bivectors $P_{A}$ and $P_{B}$ is defined by the following almost Poisson brackets between coordinates
$\{x_{i},x_{j}\}_{A}=\sum_{i<j,\,i,j\neq k}c^{k}_{ij}x_{k}\,,\qquad\{x_{i},x_{j}\}_{B}=\sum_{i<j,\,i,j\neq k}c^{k}_{ij}\lambda_{k}x_{k}$
which does not satisfy the Jacobi identity in the general case.
In modern terms Volterra proposed formal Hamiltonian description (1.14) for the second order homogeneous vector field (1.12). Using the Jacobi last multiplier theory he also proved that the phase flow of the formal Hamiltonian system (1.14) preserves the volume $n$-form
$dx_{1}\wedge\cdots\wedge dx_{n}$
if all the numbers $\lambda_{1},\cdots,\lambda_{n}$ are pairwise different. Discussion of the Volterra formal Hamiltonian description (1.14) may be found in [17, 18].
Some examples of differential equations of the form (1.14) in three-dimensional space were considered by Volterra in connection with studying the dynamics of the Earth’s poles, see collection of papers [16]. Recently these Volterra calculation was repeated for other dynamical systems in three-dimensional space [19].
@@ 1.2 Invariant Volterra brackets
$V=\sum_{i=1}^{n}V_{i}(x_{1},\ldots,x_{n})\frac{\partial}{\partial x_{i}}$
with the homogeneous components
$V_{i}(\lambda x_{1},\ldots,\lambda x_{n})=\lambda^{\alpha}V_{i}(x_{1},\ldots,x_{n})\,,\qquad i=1,\ldots,n\,.$
The Lie bracket between $V$ and the Euler vector field $U$ is equal to
$V,U]=(\alpha-1)V\,,\qquad U=\sum_{i=1}^{n}x_{i}\frac{\partial}{\partial x_{i}}\,.$
according to the Euler theorem
$\sum_{k=1}^{n}x_{k}\frac{\partial V_{i}(x_{1},\ldots,x_{n})}{\partial x_{k}}=\alpha V_{i}(x_{1},\ldots,x_{n})\,,\qquad i=1,\ldots,n.$ (1.15)
Bivector
$P^{\prime}=V\wedge U$ (1.16)
is a tensor invariant of the phase flow associated with $V$ since
$\mathcal{L}_{V}\,P^{\prime}=[V,V]\wedge U+V\wedge[V,U]=0\,.$
Here $\mathcal{L}_{V}\,P^{\prime}$ is a Lie derivative of the tensor field $P^{\prime}$ along the vector field $V$. Moreover, bivector $P^{\prime}$ (1.16) satisfies to the Jacobi condition since
$P^{\prime},P^{\prime}]\!]=2V\wedge[U,V]\wedge U=0\,$,
where [[.,.]][\![.,.]\!] is a Schouten-Nijenhuis bracket.
Let us suppose that function $F(x)$ is a homogeneous first integral of the vector field $V$, i.e.
$V(F)=\sum_{i=1}^{n}V_{i}\frac{\partial F}{\partial x_{i}}=0\qquad\mbox{and}\qquad U(F)=\sum_{i=1}^{n}x_{i}\frac{\partial F}{\partial x_{i}}=\kappa F\,.$
@@
$P^{\prime}dF=-V(F)\,U+U(F)\,V=\kappa F\,V\,$,
that allows us to represent homogeneous vector field $V$ in the formal Hamiltonian form
$V=P^{\prime}dH\,,\qquad H=\kappa^{-1}\ln|F|\,.$ (1.17)
Here $P^{\prime}$ is the rank two Poisson bivector (1.16) and $H$ is a Hamilton function.
We suppose that if $V=PdH$ is a Hamiltonian vector field defined by homogeneous $P$ and $H$, the sum
$\tilde{P}=P^{\prime}-\alpha(k,\kappa)HP$
is an invariant Poisson bivector at the special value of $\alpha(k,\kappa)$, rank $\tilde{P}$=rank $P$, i.e.
$\tilde{P},\tilde{P}]\!]=2\alpha[\![P^{\prime},HP]\!]+\alpha^{2}[\![HP,HP]\!]=2\alpha[\![PdH\wedge U,HP]\!]+\alpha^{2}[\![HP,HP]\!]=0\,.$
As an example let us consider canonical Lie-Poisson bracket on the algebra $so^{*}=so^{*}(2)\oplus so^{*}(2)$
$\{s_{i},s_{j}\}=\varepsilon_{ijk}s_{k}\qquad\mbox{and}\qquad\{t_{i},t_{j}\}=\varepsilon_{ijk}t_{k}\,,\qquad i,j,k=1,2,3,$ (1.18)
and the corresponding homogeneous bivector $P_{s}$. Here $\varepsilon_{ijk}$ is the Levi-Civita symbol.
@@
Associated with homogeneous quadratic Hamiltonian
$H=\sum_{i,j=1}^{6}a_{ij}x_{i}x_{j},\qquad x=(s_{1},s_{2},s_{3},t_{1},t_{2},t_{3})$
and homogeneous bivector $P_{s}$ (1.18) vector field $V=P_{s}dH$ together with the Euler vector field
$U=\sum x_{i}\frac{\partial}{\partial x_{i}}$
define rank two invariant Poisson bivector
$P^{\prime}=V\wedge U\,$,
whereas invariant Poisson bivector
$\tilde{P}=P^{\prime}-HP$
has a rank four.
The proof consists of a direct verification.
As a second example let us consider canonical Lie-Poisson bracket on the Lie-Poisson algebra $e^{*}(3)=so^{*}(3)\ltimes{\mathbb{R}^{3}}^{*}$
$\{L_{i},L_{j}\}=\varepsilon_{ijk}L_{k},\qquad\{L_{i},p_{j}\}=\varepsilon_{ijk}p_{k},\qquad\{p_{i},p_{j}\}=0\,$ (1.19)
and the corresponding homogeneous bivector $P_{e}$. For homogeneous quadratic Hamiltonian $H$ on $e^{*}(3)$ we have an analog of the previous Proposition which can be extended for the case of quasi-homogeneous Hamiltonian.
@@
Associated with quasi-homogeneous Hamiltonian
$H=\sum_{i,j=1}^{2}a_{ij}L_{i}L_{j}+b_{1}p_{1}+b_{2}p_{2}+b_{3}p_{3},$
and bivector $P_{e}$ (1.19) vector field $V=P_{e}dH$ together with the Euler-type vector field
$U^{\prime}=\frac{1}{2}\sum_{i=1}^{3}L_{i}\frac{\partial}{\partial L_{i}}+\sum_{i=1}^{3}p_{i}\frac{\partial}{\partial p_{i}}$
define rank two invariant Poisson bivector
$P^{\prime}=V\wedge U^{\prime}\,$,
whereas invariant Poisson bivector
$\tilde{P}=P^{\prime}-\frac{1}{2}HP$
has a rank four.
The proof consists of a direct verification.
2 Homogeneous potentials on $T^{*}\mathbb{R}^{n}$
Let us define a tensor field $P^{\prime}$ of valence $(2,0)$ that depends on the function $V(q)$ as follows
$P^{\prime}=$ $\left(\frac{p_{i}q_{j}}{m_{i}}-\frac{p_{j}q_{i}}{m_{j}}\right)\frac{\partial}{\partial q_{i}}\wedge\frac{\partial}{\partial q_{j}}+\left(\frac{k}{2m_{i}}p_{i}p_{j}+q_{i}\frac{\partial V}{\partial q_{j}}\right)\frac{\partial}{\partial q_{i}}\wedge\frac{\partial}{\partial p_{j}}$
$+$ $\frac{k}{2}\left(p_{i}\frac{\partial V}{\partial q_{j}}-p_{j}\frac{\partial V}{\partial q_{i}}\right)\frac{\partial}{\partial p_{i}}\wedge\frac{\partial}{\partial p_{j}}\,.$ (2.1)
According to this definition, the tensor field $P^{\prime}$ (2.1) is a skew-symmetric field, i.e., a bivector.
Proposition 3
If $V(q)$ is a homogeneous function of degree $k$, then the tensor field $P^{\prime}$ (2.1) is a Poisson bivector that is invariant with respect to the phase flow generated by the Hamiltonian vector field $X$ (1.4).
Proof: The derivative of $P^{\prime}$ (2.1) along the vector field $X$ (1.4) is equal to
$\mathcal{L}_{X}P^{\prime}=\frac{p_{i}}{m_{i}}\frac{\partial Z}{\partial q_{j}}\frac{\partial}{\partial q_{i}}\wedge\frac{\partial}{\partial p_{j}}-\left(\frac{\partial Z}{\partial q_{i}}\frac{\partial V}{\partial q_{j}}-\frac{\partial Z}{\partial q_{j}}\frac{\partial V}{\partial q_{i}}\right)\frac{\partial}{\partial p_{i}}\wedge\frac{\partial}{\partial p_{j}}\,$,
and, therefore, for homogeneous potentials $V(q)$, the Lie derivative is equal to zero
$\mathcal{L}_{X}P^{\prime}=0\,$,
since $Z=0$ according to Euler’s theorem (1.3).
The verification of Jacobi’s identity (1.10) for arbitrary smooth functions $f,g,$ and $h$ of variables $x=(q,p)$
$\{\{f,g\},h\}+\{\{g,h\},f\}+\{\{h,f\},g\}=0$
is reduced to checking Jacobi’s conditions for the Poisson brackets between variables $x=(q,p)$, which have the form
$\{q_{i},q_{j}\}^{\prime}=$ $\frac{p_{i}q_{j}}{m_{i}}-\frac{p_{j}q_{i}}{m_{j}}\,,\qquad\{q_{i},p_{j}\}^{\prime}=\frac{k}{2m_{i}}p_{i}p_{j}+q_{i}\frac{\partial V}{\partial q_{j}}$
$\{p_{i},p_{j}\}^{\prime}=$ $\frac{k}{2}\left(p_{i}\frac{\partial V}{\partial q_{j}}-p_{j}\frac{\partial V}{\partial q_{i}}\right)\,.$ (2.2)
For example, let’s write out one of these Jacobi’s conditions
$\{\{q_{i},q_{j}\},p_{k}\}+\{\{q_{j},p_{k}\},q_{i}\}+\{\{p_{k},q_{i}\},q_{j}\}=-\left(\frac{p_{i}q_{j}}{m_{i}}-\frac{p_{j}q_{i}}{m_{j}}\right)\dfrac{\partial}{\partial q_{k}}Z=0\,$,
where $i,j,k=1,\ldots,n$. This is true because $Z=0$ according to Euler’s theorem (1.3). Similarly, Jacobi’s conditions for the Poisson brackets between the remaining coordinates $x=(q,p)$ are also verified.
When $k\neq-2$, the invariant Poisson bivector $P^{\prime}$ (2.1) does not compatible with the Poisson bivector P{P} (1.6), i.e.
$P,\hat{P}]\!]\neq 0.$
In the case of a potential Jacobian field at k=2k=-2, the compatibility condition is satisfied $P,\hat{P}]\!]=0$ and there is another scalar invariant
$F=\frac{1}{2}\sum_{i,j=1}^{n}(q_{i}p_{j}-q_{j}p_{i})^{2}+2V\sum_{i=1}^{n}q_{i}^{2}\,,\qquad\mathcal{L}_{X}F=0\,.$
[…]
$P^{\prime}dH=mHX\,.$
According to [20] bivector $P^{\prime}$ is a product of invariant vector field $X=PdH$ and non-invariant Euler-type vector field $Y$ [21]
$P^{\prime}=X\wedge Y\,,\qquad Y=\sum_{i=1}^{n}q_{i}\frac{\partial}{\partial q_{i}}+\frac{k}{2}\sum_{i=1}^{n}p_{i}\frac{\partial}{\partial p_{i}}\,$,
so that
$\mathcal{L}_{X}Y\wedge X=0\,.$
Proposition 4
If $V(q)$ is a homogeneous function of degree $k\neq\pm 2$, then the tensor field of valence $(0,2)$
@@ 2 Homogeneous potentials on $T^{*}\mathbb{R}^{n}$
is a non-degenerate Poisson bivector, and the tensor field of valence $(2,0)$
@@
is an invariant simplectic form with respect to the phase flow generated by the Hamiltonian vector field $X$ (1.4).
The proof consists of a direct verification of non-degeneracy, invariance, and Jacobi identities
@@
which is based on the fulfillment of the homogeneity condition $Z=0$ (1.3). The invariant symplectic 2-form is equal to
$\hat{\omega}=\frac{2}{(k-2)H}\left(\frac{k}{(k+2)H}dH\wedge d\alpha^{\prime}+\sum_{i=1}^{n}dq_{i}\wedge dp_{i}\right)\,$,
where 1-form $d\alpha^{\prime}$
$d\alpha^{\prime}=\sum_{i=1}^{n}p_{i}dq_{i}-\frac{2}{k}\sum_{i=1}^{n}q_{i}dp_{i}$
is the standard Liouville form $\sum_{i=1}^{n}p_{i}dq_{i}$ on the cotangent bundle $T^{*}\mathbb{R}^{n}$ (or Poincaré 1-form or canonical form or tautological form) minus 1-form $\frac{2}{k}\sum_{i=1}^{n}q_{i}dp_{i}$.
According to the Lie-Darboux theorem, there exists a canonical transformation
$(q^,p^)\phi$$\to(\hat{q},\hat{p})$
such that
@@
at least locally. This transformation can be viewed as analogous to the Bäcklund transformations for the integrable Hamiltonian systems with homogeneous potentials [22].
The following equation
$\hat{P}dH=\frac{k+2}{2}H\,PdH$
can be considered as a tensor analogue of the Lagrange identity (1.2).
It should be noted that within the framework of Galois theory of differential groups for Hamiltonian systems with homogeneous potentials [23], cases with homogeneity index $k=\pm 2$ are also of particular significance and require separate consideration.
3 Second family of potentials obtained
Using the shift $q_{i}\to q_{i}+a_{i}$ and the appropriate limit for the parameters $k$ and $a_{i}$, we obtain the equation for the new potential $U(q)$
$\bar{Z}=\sum_{i=1}^{n}a_{i}\frac{\partial}{\partial q_{i}}\,U(q_{1},\ldots,q_{n})-kU(q_{1},\ldots,q_{n})=0\,$, (3.1)
instead of Euler’s equation (1.3) and the condition

### v4

@@ 1 Introduction
$H=T(p)+V(q)=\frac{1}{2m_{1}}p_{1}^{2}+\cdots+\frac{1}{2m_{n}}p_{n}^{2}+V(q)\,,\qquad m_{i}>0\,$, (1.2)
is the sum of kinetic energy $T(p)$ and a smooth homogeneous function $V(q)$ of degree $k$, then Lagrange’s identity holds
@@
Jacobi used this identity considering the stability of the motion of systems of $n$ interacting particles with masses $m_{1},\ldots,m_{n}$ arising in celestial mechanics [1], see also [2]. Various generalizations and applications of Lagrange’s identity are discussed in the works of V.V. Kozlov [3, 4, 5].
@@
In this note we prove that if potential $V(q)$ is a smooth homogeneous function of degree $k$, then for the Hamiltonian system (1.1,1.6) not only Lagrange’s identity (1.3) holds, but also exist an invariant Poisson bivector $\hat{P}$ and an invariant symplectic form $\hat{\omega}=\hat{P}^{-1}$, which cannot be obtained by differentiation and tensor operations from the basic tensor invariants $X,H,P$ and $\omega$.
@@ 1.1 Basic tensor invariants of Hamiltonian systems
$\Omega=\wedge^{n}\omega=(-1)^{n(n+1)/2}\,n!\,dq_{1}\wedge\cdots\wedge dq_{n}\wedge dp_{1}\cdots\wedge dp_{n}\,,\qquad\mathcal{L}_{X}\Omega=0\,.$
@@
1.2 Homogeneous vector fields
@@
$X=\sum_{i=1}^{n}X_{i}(x_{1},\ldots,x_{n})\frac{\partial}{\partial x_{i}}$
whose elements $X_{i}(x_{1},\ldots,x_{n})$ are homogeneous functions of the variables $x_{1},\ldots,x_{n}$ with homogeneity index $\alpha$
$X_{i}(\lambda x_{1},\ldots,\lambda x_{n})=\lambda^{\alpha}X_{i}(x_{1},\ldots,x_{n})\,,\qquad i=1,\ldots,n\,.$
According to Euler’s theorem on differentiable homogeneous functions, the Lie bracket between the vector field $X$ and the Euler vector field $E$ is equal to
$E,X]=(1-\alpha)X\,,\qquad E=\sum_{i=1}^{n}x_{i}\frac{\partial}{\partial x_{i}}\,.$
It follows, therefore, that the bivector
$P_{e}=X\wedge E$ (1.13)
is a tensor invariant of the phase flow defined by the vector field $X$
$\mathcal{L}_{X}\,P_{e}=[X,X]\wedge E+X\wedge[X,E]=0\,.$
which satisfies to the Jacobi condition
$P_{e},P_{e}]\!]=2X\wedge[E,X]\wedge E=0\,.$
Suppose that the function $F(x_{1},\ldots,x_{n})$ is a homogeneous first integral of the vector field $X$ with homogeneity index $\kappa$, i.e.
$X(F)=\sum_{i=1}^{n}X_{i}\frac{\partial F}{\partial x_{i}}=0\qquad\mbox{and}\qquad E(F)=\sum_{i=1}^{n}x_{i}\frac{\partial F}{\partial x_{i}}=\kappa F\,.$
@@
$P_{e}dF=-X(F)\,E+E(F)\,X=\kappa F\,X\,$,
that allows us to represent a homogeneous vector field $X$ in formal Hamiltonian form
$X=P_{e}dH\,,\qquad H=\kappa^{-1}\ln|F|\,.$
Here, $P_{e}$ is the second-rank Poisson bivector (1.13), and $H$ is the Hamiltonian.
According to [18], we call the vector field $X=P_{e}dH$ a formal Hamiltonian field because the Poisson bivector $P_{e}$ (1.13) does not possess m=n2m=n-2 uniquely determined and independent Casimir functions $C_{1},\ ...,C_{m}$ such that $P_{e}dC_{i}=0$, $i=1,\ldots,m$.
2 Homogeneous potentials
In the general case, the vector field $X$ (1.5) arising in Hamiltonian mechanics is not homogeneous. However, if the potential $V(q)$ is a smooth homogeneous function, then the Hamiltonian $H=T+V$ (1.2) consists of two homogeneous functions $T$ and $V$, which allows us to construct an analogue of the invariant Poisson bivector (1.13).
Indeed, let us define an Euler-type vector field
$Y=\sum_{i=1}^{n}q_{i}\frac{\partial}{\partial q_{i}}+\frac{k}{2}\sum_{i=1}^{n}p_{i}\frac{\partial}{\partial p_{i}}\,$, (2.1)
for which the sum of two homogeneous functions $H=T+V$ (1.2) will be a Darboux invariant [17]
$\mathcal{L}_{Y}H\equiv Y(H)=kH$
provided that the condition $Z=0$ (1.4) holds. The corresponding vector field $X=PdH$ (1.5) will also be a Darboux invariant of the vector field $Y$
$\mathcal{L}_{Y}X\equiv[Y,X]=\left(\frac{k}{2}-1\right)X+PdZ\,,\qquad\mbox{where}\qquad Z=0\,.$ (2.2)
Analogous to the bivector $P_{e}=X\wedge E$ (1.13) considered in the previous section, the bivector
$P^{\prime}=X\wedge Y\,$, (2.3)
is invariant with respect to the flow of the vector field $X$
$\mathcal{L}_{X}P^{\prime}=[X,X]\wedge Y+X\wedge[X,Y]=0$
and satisfies the Jacobi condition
$P^{\prime},P^{\prime}]\!]=2X\wedge[Y,X]\wedge Y=0\,.$
which is a direct consequence of equation (2.2), valid only when the Lagrange identity (1.3) holds.
We have thus proven the following statement.
@@ 2 Homogeneous potentials
If the Hamiltonian function $H=T(p)+V(q)$ (1.2) is a Darboux invariant of the vector field $Y$ (2.1)
$Y(H)=kH\qquad\mbox{at}\qquad H\neq 0\,$,
then the tensor field $P^{\prime}$ (2.3) is a Poisson bivector invariant with respect to the phase flow generated by the Hamiltonian vector field $X$ (1.5).
In terms of the variables $q$ and $p$, this Poisson bivector takes the following form
$P^{\prime}=$ $\sum_{1\leq i<j\leq n}\left(\frac{p_{i}q_{j}}{m_{i}}-\frac{p_{j}q_{i}}{m_{j}}\right)\frac{\partial}{\partial q_{i}}\wedge\frac{\partial}{\partial q_{j}}+\sum_{i=1}^{n}\sum_{j=1}^{n}\left(\frac{k}{2m_{i}}p_{i}p_{j}+q_{i}\frac{\partial V}{\partial q_{j}}\right)\frac{\partial}{\partial q_{i}}\wedge\frac{\partial}{\partial p_{j}}$
$+$ $\sum_{1\leq i<j\leq n}\frac{k}{2}\left(p_{i}\frac{\partial V}{\partial q_{j}}-p_{j}\frac{\partial V}{\partial q_{i}}\right)\frac{\partial}{\partial p_{i}}\wedge\frac{\partial}{\partial p_{j}}\,.$
For $k\neq-2$, the invariant Poisson bivector $P^{\prime}$ (2.3) does not compatible with the Poisson bivector $P$ (1.7) since
$P,P^{\prime}]\!]\neq 0.$
Recall that the Poisson bivectors $P$ and $P^{\prime}$ are compatible if any linear combination $(aP+bP^{\prime})$ with constant coefficients $a$ and $b$ is a Poisson tensor, that is, the Jacobian identity holds
$aP+bP^{\prime},aP+bP^{\prime}]\!]=2ab[\![P,P^{\prime}]\!]=0\,,\qquad\forall a,b\in\mathbb{C}.$
In the case of the Jacobi potential field for k=2k=-2, the compatibility condition holds: $P,{P}^{\prime}]\!]=0$, and in this case there is another scalar invariant
$F=\frac{1}{2}\sum_{i,j=1}^{n}(q_{i}p_{j}-q_{j}p_{i})^{2}+2V\sum_{i=1}^{n}q_{i}^{2}\,,\qquad\mathcal{L}_{X}F=0\,$,
see the definition of the Jacobi potential field and the corresponding discussion in [5].
The Poisson bivector $P^{\prime}$ (2.3) is degenerate for $n\geq 2$, since $P^{\prime}$ has rank 2, and satisfies the relation
$P^{\prime}dH=kHX\,.$
In the general case, the Poisson bivector $P^{\prime}$ (2.3) does not compatible with the Poisson bivector P{P} (1.7). Nevertheless, there exists a unique linear combination of these invariant bivectors with coefficients depending on the Hamiltonian function $H$ and the degree of homogeneity of the potential $k$, which is an invariant Poisson tensor.
@@
If $V(q)$ is a homogeneous function of degree $k\neq\pm 2$, then on the open subset defined by the inequality $H\neq 0$, the tensor field of valence $(2,0)$
@@
is a non-degenerate Poisson bivector. This allows us to define on this open subset a tensor field of valence $(0,2)$
@@
which is a symplectic form invariant with respect to the phase flow generated by the Hamiltonian vector field $X$ (1.5).
The proof consists of directly verifying the non-degeneracy, invariance, and the Jacobi identity
@@
which are equivalent to the condition $Z=0$ (1.4).
According to Lie-Darboux theorem, there exists a coordinate transformation $(q^,p^)\phi$$\to(\hat{q},\hat{p})$ such that
@@
at least locally. This transformation can be regarded as an analogue of the Baecklund transformations for integrable Hamiltonian systems with homogeneous potentials [19].
For all three invariant Poisson bivectors $P$, $P^{\prime}$, and $\hat{P}$, the Hamiltonian flows generated by the Hamiltonian function $H$ with potential $V(q)$ (1.4) are proportional to each other
$P^{\prime}dH=kHX\qquad\mbox{and}\qquad\hat{P}dH=\frac{k+2}{2}H\,X$
where $X=PdH$.
We note that the cases $k=\pm 2$ are special cases within the application of Galois group theory to Hamiltonian systems with homogeneous potentials [20].
3 Nonhomogeneous potentials of a special kind
Let us consider the Hamilton function
$H=\frac{1}{2m_{1}}p_{1}^{2}+\cdots+\frac{1}{2m_{n}}p_{n}^{2}+U(q)$
where the potential $U(q)$ satisfies the relation
$\bar{Z}=0\,,\qquad\mbox{where}\qquad\bar{Z}=\sum_{i=1}^{n}a_{i}\frac{\partial}{\partial q_{i}}\,U(q_{1},\ldots,q_{n})-kU(q_{1},\ldots,q_{n})\,.$ (3.1)
In this case, the equivalent of Lagrange’s identity (1.3) is the expression